When a skydiver jumps out of an airplane she initially accelerates towards the earth at 9.81m/s squared. Wind resistance increases on the skydiver as they fall faster and faster through the air and eventually the skydiver reaches terminal velocity at 55.6m/s. When the diver opens her parachute, she experiences an acceleration of 30.0m/s squared. If the velocity must be 5.0 m/s at landing to avoid serious injury, then the minimum height above the ground that she must open her parachute would be ? __________ M
1
Expert's answer
2020-11-30T14:53:19-0500

"v=v_0-at\implies t= \frac{v_0-v}{a},"

"h=v_0t-\frac{at^2}{2}=\frac{v_0^2-v^2}{2a}=\frac{55.6^2-5^2}{2 \cdot 30}=51 ~m."